OpenStudy (anonymous):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. A. -2 and -4 B. -2 and 4 C. 2 and -4 D. 2 and 4

zepdrix (zepdrix):

Do you understand how to translate the sentences into a system of equations? :)

OpenStudy (anonymous):

no

zepdrix (zepdrix):

The sum of 5 times `a number` and twice `a second number` is 18 Translates to: 5x + 2y = 18 So I have 5 times something, I'm calling that something x. and 2 times something else, calling that something else y.

zepdrix (zepdrix):

5x+2y=18

zepdrix (zepdrix):

The difference of three times `the first number` and seven times `the second number` is -22 How bout this second sentence, any ideas? :)

OpenStudy (anonymous):

3x+7y=-22 that's probably wrong

zepdrix (zepdrix):

Mmm very close! ^^ When they use the word `difference`, they want us to subtract the two terms, not add them.

zepdrix (zepdrix):

So, something like that, yah? 3x-7y=-22

OpenStudy (anonymous):

oh okay but im still confused on the question because they don't have 2 or 4 in their sentence

zepdrix (zepdrix):

So we've established a system of equations:\[\Large\rm 5x+2y=18\]\[\Large\rm 3x-7y=-22\]Two equations, two unknowns, so we can solve this. We want to solve this using either `substitution` or `elimination`. It looks like `elimination` will work a lot better for this particular problem. You are familiar with that method? :)

zepdrix (zepdrix):

For this problem, let's try to match up our y's. We want the same coefficient on both y's.

zepdrix (zepdrix):

So we need the LCM (Least common multiple) of 2 and 7.

zepdrix (zepdrix):

Which just ends up being 14 in this case. (2 x 7). So we want to multiply the first equation (the whole thing) by 7, while we'll multiply the second equation by 2, to get our 14s

zepdrix (zepdrix):

Multiplying everything by 7 in the first equation gives us: \[\Large\rm 7(5x+2y=18)\]\[\Large\rm 35x+14y=126\] And for our second equation:\[\Large\rm 2(3x-7y=-22)\]\[\Large\rm 6x-14y=-88\]I multiplied everything by 2.

OpenStudy (anonymous):

okay so I tried to do that to the numbers in the question I did 18 - -22 and got -4

zepdrix (zepdrix):

Well we had to do this work getting our y's matched up before we could add or subtract the equations. \[\Large\rm 35x+\color{orangered}{14y}=126\]\[\Large\rm 6x-\color{orangered}{14y}=-88\]

zepdrix (zepdrix):

If we ADD these equations now, the y's will cancel out.

zepdrix (zepdrix):

35 + 6 gives us 41 x's.

OpenStudy (anonymous):

oh ok

zepdrix (zepdrix):